THe length of a rectangular pool is 6 meters less than twice it's width. If the pool's perimeter is 126 meters, what are its dimensions?
I came up with the equation 6L-2w= 126. I just don't know where to go from there.
Thank you

Answer by checkley77(12844)

L=2W
2L+2W=126
2(2W)+2W=126
4W+2W=126
6W=126
W=21 METERS IS THE WIDTH.
L=2*21=42 METERS IS THE LENGTH.
PROOF:
2*42+2*21=126
84+42=126
126=126